# Simple-RF: Regularizing Sparse Input Radiance Fields with Simpler Solutions

### v2

@@ 1. Introduction
We design different augmentations for NeRF, TensoRF and ZipNeRF based on different shortcomings and architectures of these models. The high positional encoding degree used in the NeRF leads to undesired depth discontinuities, creating floaters. Further, the view-dependent radiance feature leads to shape-radiance ambiguity, creating duplication artifacts. We design augmentations for the NeRF by reducing the positional encoding degree and disabling the view-dependent radiance feature. On the other hand, the large number of high-resolution factorized components in TensoRF and the large hash table in ZipNeRF cause floaters in these models in the few-shot setting. Thus, we design augmentations to constrain the model with respect to such components to learn simpler solutions.
@@
  • We show that our regularization achieves substantial improvements on different radiance fields and on four different datasets.
@@ 2.1. Implicit Radiance Fields
Deep Learning Based Depth Priors: There exist multiple models that utilize the advances in dense depth estimation using deep neural networks. DDP-NeRF (Roessle et al., 2022) extends DS-NeRF by employing a CNN to complete the sparse depth into dense depth for more supervision. SCADE (Uy et al., 2023) and SparseNeRF (Wang et al., 2023) use the depth map output by single image depth models to constrain the absolute and the relative order of pixel depths, respectively. DiffusioNeRF (Wynn and Turmukhambetov, 2023) learns the joint distribution of RGBD patches using denoising diffusion models (DDM) and utilizes the gradient of the distribution provided by the DDM to regularize NeRF rendered RGBD patches. However, the deep-learning based priors require pre-training on a large dataset and may suffer from generalization issues when obtaining the prior on unseen test scenes. Our work obtains depth supervision by harnessing the power of learning without suffering from generalization issues by employing and training augmented models on the given scene alone.
@@ 4. Method
Learning a radiance field with sparse input views leads to overfitting on the input views with severe distortions in novel views. Our key observation is that most of the distortions are due to the sub-optimal use of the high capabilities of the radiance field model. Further, we find that reducing the capability of the radiance field helps constrain the model to learn only simpler solutions, which can provide better depth supervision in certain regions of the scene. However, the lower capability models are not optimal either since they cannot learn complex solutions where necessary. Our solution here is to use the higher capability model as the main model and employ the lower capability models as augmentations to provide guidance on where to use simpler solutions. The challenge is that, it is not known apriori where one needs to employ supervision from the augmented model. We determine the more accurate model among the main and augmented models in terms of the estimated depth and use the more reliable depth to supervise the other. We note that the augmented models are employed only during the learning phase and not during inference. Thus, there is no additional overhead during inference. The augmented models are similar to the main model, but we modify their parameters to reduce their capability, and train them in tandem with the main model.
@@ 4.2.3. Design of Augmentations
where $H(w_{1},w_{2},\ldots,w_{n})=\sum_{i=1}^{n}(-w_{i}\log w_{i})$ is the entropy operator, $N$ is the number of 3D points $\mathbf{p}_{i}$ along a ray and $w_{i}$ is the weight corresponding to $\mathbf{p}_{i}$ as described in Eq. 2.
@@ 5.1. Datasets
Table 2. Quantitative results of NeRF based models on the NeRF-LLFF dataset.
@@
Table 3. Quantitative results of NeRF based models on the RealEstate-10K dataset.
@@ 5.2. Evaluation measures
We quantitatively evaluate the predicted frames from various models using the peak signal-to-noise ratio (PSNR), structural similarity (SSIM) (Wang et al., 2004) and LPIPS (Zhang et al., 2018) measures. For LPIPS, we use the v0.1 release with the AlexNet (Krizhevsky et al., 2012) backbone, as suggested by the authors. We also employ depth mean absolute error (MAE) to evaluate the models on their ability to predict absolute depth in novel views. In addition, we also evaluate the models with regard to their ability to predict better relative depth using the spearman rank order correlation coefficient (SROCC). Obtaining better relative depth might be more crucial in downstream applications such as 3D scene editing. Since the ground truth depth is not provided in the datasets, we train NeRF and ZipNeRF models with dense input views on forward-facing and 360 datasets respectively and use their depth predictions as pseudo ground truth. On the NeRF-LLFF and MipNeRF360 datasets, we normalize the predicted depths by the median ground truth depth, since the scenes have different depth ranges. With very few input views on forward-facing datasets, the test views could contain regions that are not visible in the input views, and hence, we also evaluate both the view synthesis and depth performance in visible regions only. To determine such regions, we use the depth estimated by a NeRF trained with dense input views and compute the visible region mask through reprojection error in depth. We provide more details on the mask computation in the supplementary. On the other hand, the input views cover most of the scene in the 360 datasets, and hence we evaluate the performance on full frames. We do not evaluate the rendered depth on the NeRF-Synthetic dataset since the depth estimated with the dense input ZipNeRF is unreliable, especially in the white background regions, and the ground truth depth is not provided in the dataset either.
@@ 6.1. Simple-NeRF
Figure 10. Qualitative examples of NeRF based models on RealEstate-10K dataset with three input views. Simple-NeRF predictions are closest to the ground truth among all the models. In particular, DDP-NeRF predictions have a different shade of color and ViP-NeRF suffers from shape-radiance ambiguity, creating ghosting artifacts.
Figure 11. Qualitative examples of NeRF based models on the RealEstate-10K dataset with four input views. We find that Simple-NeRF and ViP-NeRF perform the best among all the models. However, ViP-NeRF predictions contain minor distortions, as pointed out by the magenta arrow, which is rectified by Simple-NeRF.
Figure 12. Qualitative examples of NeRF based models on the NeRF-LLFF dataset with two input views. DDP-NeRF and ViP-NeRF synthesize frames with broken objects in the second row, and FreeNeRF breaks the object in the first row due to incorrect depth estimations. Simple-NeRF produces sharper frames devoid of such artifacts.
Figure 13. Qualitative examples of NeRF based models on the NeRF-LLFF dataset with three input views. In the first row, the orchid is displaced out of the cropped box in the FreeNeRF prediction, due to incorrect depth estimation. ViP-NeRF and RegNeRF fail to predict the complete orchid accurately and contain distortions at either end. In the second row, ViP-NeRF prediction contains severe distortions. Simple-NeRF reconstructs the best among all the models in both examples.
Figure 14. Qualitative examples of NeRF based models on the NeRF-LLFF dataset with four input views. In the first row, we find that ViP-NeRF, FreeNeRF, and DDP-NeRF struggle to reconstruct the shape of the leaf accurately. In the second row, DS-NeRF introduces floaters. Simple-NeRF does not suffer from such artifacts and reconstructs the shapes better.
@@ 6.1.1. Comparisons
We evaluate the performance of our Simple-NeRF on the two forward-facing datasets only since the NeRF does not natively support unbounded 360 scenes. We evaluate the performance of our model against various sparse input NeRF models. We compare with DS-NeRF (Deng et al., 2022), DDP-NeRF (Roessle et al., 2022) and RegNeRF (Niemeyer et al., 2022) which regularize the depth estimated by the NeRF. We also evaluate DietNeRF (Jain et al., 2021) and InfoNeRF (Kim et al., 2022) that regularize the NeRF in hallucinated viewpoints. We also include two recent models, FreeNeRF (Yang et al., 2023) and ViP-NeRF (Somraj and Soundararajan, 2023), among the comparisons. We train the models on both datasets using the codes provided by the respective authors.
@@ 6.1.3. Quantitative and Qualitative Results
Tabs. 2 and 3 show the view-synthesis performance of Simple-NeRF and other prior art on NeRF-LLFF and RealEstate-10K datasets respectively. We find that Simple-NeRF achieves state-of-the-art performance on both datasets in most cases. The higher performance of all the models on the RealEstate-10K dataset is perhaps due to the scenes being simpler. Hence, the performance improvement is also smaller as compared to the NeRF-LLFF dataset. Fig. 9 shows predictions of various models on an example scene from the RealEstate-10K dataset, where we observe that Simple-NeRF is the best in reconstructing the novel view. Figs. 10, 11, 12, 13, 14 and 15 show more comparisons on both datasets with 2, 3, and 4 input views. Further, Simple-NeRF improves significantly in estimating the depth of the scene as seen in  Tab. 4 and  Fig. 16. We provide video comparisons in the supplementary.
@@ 6.1.4. Ablations
(b) Without Lambertian augmentation: The ablated model suffers from shape-radiance ambiguity and produces ghosting artifacts.
@@ 6.3. Simple-ZipNeRF
Augmented ZipNeRF 0.6825 0.4462 16.27 96.42 0.49 0.619 0.5244 19.31 0.5917 0.5646 21.21
Simple-ZipNeRF 0.4876 0.5245 17.60 3.54 0.51 0.3421 0.6456 21.03 0.239 0.7458 24.19
@@
(c) 36 input views.
Figure 21. Qualitative examples of ZipNeRF and Simple-ZipNeRF on the MipNeRF360 dataset. In the first column, we observe that ZipNeRF places large regions of the pink mat close to the camera, occluding the bulldozer. In the second example, we observe objects being broken or placed at incorrect positions due to incorrect depth estimation, as well as translucent floaters in ZipNeRF predictions. Finally, in the third column, we observe that ZipNeRF fails to reconstruct the tree stump. In all the cases, Simple-ZipNeRF produces very good reconstructions without any floaters.
@@ 6.3.2. Quantitative and Qualitative Results
We compare the performance of ZipNeRF with and without our augmentations on the MipNeRF360 and NeRF-Synthetic datasets in Tabs. 7 and 8 respectively. We observe that including our augmentations improves performance significantly on both datasets in terms of all the evaluation measures. This observation is further supported by the qualitative examples in Figs. 21, 23 and 22, where we observe a clear improvement in the quality of the rendered novel views and depth when employing our augmentations. In addition, Tabs. 7 and 24 also shows the performance of the augmented model on the MipNeRF360 dataset. We observe a significant reduction in distortions in the renders of the augmented model; however, the same does not reflect in the quantitative evaluation due to the blur introduced by the augmented model.
@@ 7.2. Limitations and Future Work
Our approach of obtaining reliable depth supervision by learning simpler solutions is limited to the cases where the base model achieves a reasonable performance but suffers from various distortions due to overfitting. Our regularizations may not help significantly if the performance of the base model is very poor, as in the case of learning a highly complex scene with very few input views. In such cases, we can employ our regularizations on top of other sparse input radiance fields as we do for the NeRF-LLFF dataset.
@@
Finally, our approach of employing augmentations to obtain better supervision can be explored in sparse input novel view synthesis of highly specular objects (Verbin et al., 2022), refractive objects (Deng et al., 2024), low-light scenes (Mildenhall et al., 2022), blurred images (Ma et al., 2022) and high dynamic range images (Lu et al., 2024). Further, it would be interesting to explore the use of augmentations in related inverse problems such as surface reconstruction (Wang et al., 2021a; Long et al., 2022), dynamic view synthesis (Pumarola et al., 2021; Fridovich-Keil et al., 2023; Somraj et al., 2024, 2022) and style transfer (Wang et al., 2024) when only sparse input viewpoints are available.

### v4

@@ 1. Introduction
We design different augmentations for NeRF, TensoRF and ZipNeRF based on different shortcomings and architectures of these models. The high positional encoding degree used in the NeRF leads to undesired depth discontinuities, creating floaters. Further, the view-dependent radiance feature leads to shape-radiance ambiguity, creating duplication artifacts. We design augmentations for the NeRF by reducing the positional encoding degree and disabling the view-dependent radiance feature. On the other hand, the large number of high-resolution factorized components in TensoRF and the large hash table in ZipNeRF cause floaters in these models in the few-shot setting. The unifying theme across the three models is that we design augmentations to constrain the model with respect to such components to learn simpler solutions.
@@
  • We find that the high capability of various radiance field models lead to learning on unnecessary complex geometries when training with sparse input views. We design specific augmented models that learn simpler solutions to regularize the different radiance field models. In particular,
@@ 2.1. Implicit Radiance Fields
Deep Learning Based Depth Priors: There exist multiple models that utilize the advances in dense depth estimation using deep neural networks. DDP-NeRF (Roessle et al., 2022) extends DS-NeRF by employing a CNN to complete the sparse depth into dense depth for more supervision. SCADE (Uy et al., 2023) and SparseNeRF (Wang et al., 2023) use the depth map output by single image depth models to constrain the absolute and the relative order of pixel depths, respectively. DiffusioNeRF (Wynn and Turmukhambetov, 2023) learns the joint distribution of RGBD patches using denoising diffusion models (DDM) and utilizes the gradient of the distribution provided by the DDM to regularize NeRF rendered RGBD patches. However, the deep-learning based priors require pre-training on a large dataset and may suffer from generalization issues when obtaining the prior on unseen test scenes. Our work obtains depth supervision by harnessing the power of learning without suffering from generalization issues by employing and training augmented models on the given scene alone. In addition, our approach imposes depth supervision selectively i.e. only in the regions where the depth supervision is determined to be reliable.
@@ 4. Method
Learning a radiance field with sparse input views leads to overfitting on the input views with severe distortions in novel views. Our key observation is that most of the distortions are due to the sub-optimal use of the high capabilities of the radiance field model. Further, we find that reducing the capability of the radiance field helps constrain the model to learn only simpler solutions, which can provide better depth supervision in certain regions of the scene. However, the lower capability models are not optimal either, since they cannot learn complex solutions where necessary. Our solution here is to use the higher capability model as the main model and employ the lower capability models as augmentations to provide guidance on where to use simpler solutions. The challenge is that, it is not known apriori where one needs to employ supervision from the augmented model. We determine the more accurate model among the main and augmented models in terms of the estimated depth and use the more reliable depth to supervise the other. We note that the augmented models are employed only during the learning phase and not during inference. Thus, there is no additional overhead during inference. The augmented models are similar to the main model, but we modify their parameters to reduce their capability, and train them in tandem with the main model.
@@ 4.2.3. Design of Augmentations
where $H(w_{1},w_{2},\ldots,w_{n})=\sum_{i=1}^{n}(-w_{i}\log w_{i})$ is the entropy operator, $N$ is the number of 3D points $\mathbf{p}_{i}$ along a ray and $w_{i}$ is the weight corresponding to $\mathbf{p}_{i}$ as described in Eq. 2. The effect of ${\mathcal{L}_{mc}}$ can be observed in the supplementary videos.
@@ 4.4. Overall Loss
4.5. Discussion
We note that, although we design different augmented models for Simple-NeRF, Simple-TensoRF and Simple-ZipNeRF, the core idea behind the design of the augmentations is the same. In particular, all three radiance field models learn undesired depth discontinuities leading to floaters. We mitigate this by designing augmentations that learn smoother geometry through lower positional encoding degree in the NeRF, smaller grid in TensoRF and smaller hash-table in ZipNeRF. While the specifics of the smoothing augmentations differ, the mechanism of regularizing the main model using the learnt depth prior from the augmentations remains the same across the three Simple-RF models.
Further, our augmentations can be extended to regularize other radiance field models when training with limited views. For example, most of the NeRF family of models such as MipNeRF360 (Barron et al., 2022) and Ref-NeRF (Verbin et al., 2022) suffer from floater artifacts when trained with sparse input views. Our smoothing augmentation can be employed to mitigate the floaters. Further, Ref-NeRF (Verbin et al., 2022) extends the NeRF to handle scenes with shiny objects. When training with limited views, Ref-NeRF suffers from duplication artifacts similar to Fig. 4(b). Such artifacts can be mitigated using our Lambertian augmentation discussed in Sec. 4.1.3. Similarly, K-Planes (Fridovich-Keil et al., 2023) extends TensoRF to handle varying illumination in the input images, by learning a global appearance code per input image, similar to NeRF-W (Martin-Brualla et al., 2021). With sparse input views, the global appearance code might end up learning geometric information as well, leading to a lack of geometric details in novel views. An augmented model with a lower dimensional appearance code can help mitigate such artifacts.
@@ 5.1. Datasets
Table 2. Quantitative results of NeRF and 3DGS based models on the NeRF-LLFF dataset.
@@
Table 3. Quantitative results of NeRF and 3DGS based models on the RealEstate-10K dataset.
@@ 5.2. Evaluation measures
We quantitatively evaluate the predicted frames from various models using the peak signal-to-noise ratio (PSNR), structural similarity (SSIM) (Wang et al., 2004) and LPIPS (Zhang et al., 2018) measures. For LPIPS, we use the v0.1 release with the AlexNet (Krizhevsky et al., 2012) backbone, as suggested by the authors. We also employ depth mean absolute error (MAE) to evaluate the models on their ability to predict absolute depth in novel views. In addition, we also evaluate the models with regard to their ability to predict better relative depth using the spearman rank order correlation coefficient (SROCC). Obtaining better relative depth might be more crucial in downstream applications such as 3D scene editing. Since the ground truth depth is not provided in the datasets, we train NeRF and ZipNeRF models with dense input views on forward-facing and 360 datasets respectively and use their depth predictions as pseudo ground truth. On the NeRF-LLFF and MipNeRF360 datasets, we normalize the predicted depths by the median ground truth depth, since the scenes have different depth ranges. With very few input views on forward-facing datasets, the test views could contain regions that are not visible in the input views, and hence, we also evaluate both the view synthesis and depth performance in visible regions only. To determine such regions, we use the depth estimated by a NeRF trained with dense input views and compute the visible region mask through reprojection error in depth. We provide more details on the mask computation in the supplementary. On the other hand, the input views cover most of the scene in the 360 datasets, and hence we evaluate the performance on full frames. We do not evaluate the rendered depth on the NeRF-Synthetic dataset since the depth estimated with the dense input ZipNeRF is unreliable, especially in the white background regions, and the ground truth depth is not provided in the dataset either. We also do not evaluate the rendered depth of 3DGS based models since we find that their depth is inconsistent with NeRF based models.
@@ 6.1. Simple-NeRF
Figure 10. Qualitative examples of NeRF and 3DGS based models on RealEstate-10K dataset with three input views. Simple-NeRF predictions are closest to the ground truth among all the models. In particular, DDP-NeRF predictions have a different shade of color and ViP-NeRF suffers from shape-radiance ambiguity, creating duplication artifacts.
Figure 11. Qualitative examples of NeRF and 3DGS based models on the RealEstate-10K dataset with four input views. We find that Simple-NeRF and ViP-NeRF perform the best among all the models. However, ViP-NeRF predictions contain minor distortions, as pointed out by the magenta arrow, which is rectified by Simple-NeRF.
Figure 12. Qualitative examples of NeRF and 3DGS based models on the NeRF-LLFF dataset with two input views. DDP-NeRF and ViP-NeRF synthesize frames with broken objects in the second row, and FreeNeRF breaks the object in the first row due to incorrect depth estimations. Simple-NeRF produces sharper frames devoid of such artifacts.
Figure 13. Qualitative examples of NeRF and 3DGS based models on the NeRF-LLFF dataset with three input views. In the first row, the orchid is displaced out of the cropped box in the FreeNeRF prediction, due to incorrect depth estimation. While ViP-NeRF and RegNeRF fail to predict the complete orchid accurately causing distortions at either end, InstantSplat prediction contains severe distortions. In the second row, ViP-NeRF prediction contains severe distortions. Simple-NeRF reconstructs the best among all the models in both examples.
Figure 14. Qualitative examples of NeRF and 3DGS based models on the NeRF-LLFF dataset with four input views. In the first row, we find that ViP-NeRF, FreeNeRF, and DDP-NeRF struggle to reconstruct the shape of the leaf accurately, and CoR-GS output is excessively blurred. In the second row, DS-NeRF introduces floaters. Simple-NeRF does not suffer from such artifacts and reconstructs the shapes better.
@@ 6.1.1. Comparisons
We evaluate the performance of our Simple-NeRF on the two forward-facing datasets only since the NeRF does not natively support unbounded 360 scenes. We evaluate the performance of our model against various sparse input NeRF models. We compare with DS-NeRF (Deng et al., 2022), DDP-NeRF (Roessle et al., 2022) and RegNeRF (Niemeyer et al., 2022) which regularize the depth estimated by the NeRF. We also evaluate DietNeRF (Jain et al., 2021) and InfoNeRF (Kim et al., 2022) that regularize the NeRF in hallucinated viewpoints. We also include two recent models, FreeNeRF (Yang et al., 2023) and ViP-NeRF (Somraj and Soundararajan, 2023) and two 3DGS based models, InstantSplat (Fan et al., 2024) and CoR-GS (Zhang et al., 2024), among the comparisons. We find that InstantSplat performs better with known camera extrinsics and hence report the results with known camera extrinsics. We train the models on both datasets using the codes provided by the respective authors.
@@ 6.1.3. Quantitative and Qualitative Results
Tabs. 2 and 3 show the view-synthesis performance of Simple-NeRF and other prior art on NeRF-LLFF and RealEstate-10K datasets respectively. We find that Simple-NeRF achieves state-of-the-art performance on both datasets in most cases. The higher performance of all the models on the RealEstate-10K dataset is perhaps due to the scenes being simpler. Hence, the performance improvement is also smaller as compared to the NeRF-LLFF dataset. Fig. 9 shows predictions of various models on an example scene from the RealEstate-10K dataset, where we observe that Simple-NeRF is the best in reconstructing the novel view. In particular, we see that DDP-NeRF (Roessle et al., 2022), which uses depth estimated by deep neural network to constrain the NeRF, provides incorrect depth supervision leading to blurred floaters. On the other hand, RegNeRF (Niemeyer et al., 2022), which uses depth smoothness prior, fuses objects at different depths. In contrast to both these models, our approach of learning in-situ depth priors enables Simple-NeRF to mitigate both the issues. Figs. 10, 11, 12, 13, 14 and 15 show more comparisons on both datasets with 2, 3, and 4 input views. Further, Simple-NeRF improves significantly in estimating the depth of the scene as seen in  Tab. 4 and  Fig. 16. We provide video comparisons in the supplementary.
@@ 6.1.4. Ablations
(b) Without Lambertian augmentation: The ablated model suffers from shape-radiance ambiguity and produces duplication artifacts.
@@ 6.3. Simple-ZipNeRF
InstantSplat 0.4503 0.5204 17.57 0.4816 0.5545 18.02 0.6268 0.5461 16.37
CoR-GS 0.5039 0.5834 18.32 0.3500 0.7008 21.98 0.2239 0.7907 24.98
Augmented ZipNeRF 0.6825 0.4462 16.27 96.42 0.49 0.6190 0.5244 19.31 0.5917 0.5646 21.21
Simple-ZipNeRF 0.4876 0.5245 17.60 3.54 0.51 0.3421 0.6456 21.03 0.2390 0.7458 24.19
@@
Figure 21. Qualitative examples of ZipNeRF and 3DGS based models on the MipNeRF360 dataset. While the ZipNeRF and CoR-GS create large floaters in the first row, InstantSplat fails to reconstruct the bonsai in the second row. Simple-ZipNeRF is able to faithfully reconstruct the scene without such artifacts.
Figure 22. Qualitative examples of ZipNeRF and 3DGS based models on the MipNeRF360 dataset with 36 input views. We observe CoR-GS excessively smoothing the grass regions and ZipNeRF distorting the geometry of the bicycle. Simple-ZipNeRF reconstructs both the bicycle and the bench without smoothing the grass regions.
@@
(a) MipNeRF360 Kitchen scene with 12 input views.
(b) MipNeRF360 Bicycle scene with 12 input views.
Figure 27. Limitations of our augmentation. With 12 input views, we observe that ZipNeRF is able to reconstruct the bulldozer in the kitchen scene, but has severe floater artifacts. In this case, our augmentation is successful in removing the floaters. However, ZipNeRF is unable to reconstruct the bicycle and the bench in the bicycle scene, where our augmentation is unable to help.
@@ 6.3.2. Quantitative and Qualitative Results
We compare the performance of ZipNeRF with and without our augmentations on the MipNeRF360 and NeRF-Synthetic datasets in Tabs. 7 and 8 respectively. We observe that including our augmentations improves performance significantly on both datasets in terms of all the evaluation measures. This observation is further supported by the qualitative examples in Figs. 21, 22, 24 and 23, where we observe a clear improvement in the quality of the rendered novel views and depth when employing our augmentations. From Tab. 7, we observe that the performace of InstantSplat decreases with more number of input views. This is perhaps because of applying DUSt3R (Wang et al., 2024b) independently on each pair of input views, leading to accumulation of errors. While CoR-GS achieves very good performance, we find that it does not reconstruct high frequency details in the scene as seen in Fig. 22. As a result, it achieves better PSNR and SSIM scores, but poorer LPIPS scores. Further, we observe that CoR-GS also suffers from large floater artifacts as shown in Fig. 21(a). In addition, Tabs. 7 and 25 also shows the performance of the augmented model on the MipNeRF360 dataset. We observe a significant reduction in distortions in the renders of the augmented model; however, the same does not reflect in the quantitative evaluation due to the blur introduced by the augmented model.
@@ 7.2. Limitations and Future Work
Our approach of obtaining reliable depth supervision by learning simpler solutions is limited to the cases where the base model achieves a reasonable performance but suffers from various distortions due to overfitting. Our regularizations may not help significantly if the performance of the base model is very poor, as in the case of learning a highly complex scene with very few input views as shown in Fig. 27. In such cases, we can employ our regularizations on top of other sparse input radiance fields as we do for the NeRF-LLFF dataset.
@@
Finally, our approach of employing augmentations to obtain better supervision can be explored in sparse input novel view synthesis of highly specular objects (Verbin et al., 2022), refractive objects (Deng et al., 2024), low-light scenes (Mildenhall et al., 2022), blurred images (Ma et al., 2022) and high dynamic range images (Lu et al., 2024). Further, it would be interesting to explore the use of augmentations in related inverse problems such as surface reconstruction (Wang et al., 2021a; Long et al., 2022), dynamic view synthesis (Pumarola et al., 2021; Fridovich-Keil et al., 2023; Somraj et al., 2024, 2022; Choudhary et al., 2025) and style transfer (Wang et al., 2024a) when only sparse input viewpoints are available.